### v2

@@ 7.2. Characterization via two graphs
Now assume that ${\boldsymbol{\sigma}}$ is primitive and labels a path $\left(G^{L}_{n}\right)_{n\geq 0}$ in $\mathcal{G}^{L}(\mathcal{S})$ and a path $\left(G^{R}_{n}\right)_{n\geq 0}$ in $\mathcal{G}^{R}(\mathcal{S})$. The shift $X_{\boldsymbol{\sigma}}$ is minimal by primitiveness of ${\boldsymbol{\sigma}}$. Let $u\in\mathcal{L}(X_{\boldsymbol{\sigma}})$ be a factor. If $u$ is a $\sigma_{0}$-initial factor, then it is dendric since $\sigma_{0}$ is dendric. Otherwise, there exist a unique $k>0$ and a unique $v\in\mathcal{L}(X_{\boldsymbol{\sigma}}^{(k)})$ $\sigma_{k}$-initial such that $u$ is an extended image of $v$ by $\sigma_{0}\dots\sigma_{k-1}$ by Proposition 8. Let $Y$ be an Arnoux-Rauzy shift and let $Z=\sigma_{k}\cdot Y$. Since $v$ is $\sigma_{k}$-initial, $\mathcal{E}_{Z}(v)=\mathcal{E}_{X_{\boldsymbol{\sigma}}^{(k)}}(v)$. The graph $G^{L}(Y)$ (resp., $G^{R}(Y)$) is the complete (simple and monochromatic) graph on $\mathcal{A}$, thus $G^{L}_{k+1}$ (resp., $G^{R}_{k+1}$) is a covering subgraph. By Lemma 51 applied $k+1$ times, we obtain that $\sigma_{0}\dots\sigma_{k-1}\cdot Z$ is dendric thus, starting with the extension graph $\mathcal{E}_{Z}(v)$, $v$ only has dendric extended images under $\sigma_{0}\dots\sigma_{k-1}$ and $u$ is dendric. ∎
@@
By construction, on an alphabet $\mathcal{A}$ of size $d$, the graphs $\mathcal{G}^{L}(\mathcal{S})$ and $\mathcal{G}^{R}(\mathcal{S})$ have $d^{d-2}$ vertices (see sequence A000272 on OEIS) each, meaning that if we take the product of these graphs, we obtain a characterization with a unique graph $\mathcal{G}(\mathcal{S})$ with $d^{2(d-2)}$ vertices.
@@
However, it is possible to reduce the number of vertices with a clever use of permutations. Indeed, the vertices of $\mathcal{G}(\mathcal{S})$ are pairs of labeled trees and some of these pairs are equivalent, in the sense that they are equal up to a permutation of the letters. Thus, by keeping only one pair per equivalence class and adding permutations to the morphisms of the set $\mathcal{S}$, we obtain a smaller graph.
@@
7.3. An example: $\mathcal{S}$-adic characterization of minimal dendric shifts with one right special factor of each length
While we do not have an easy description of an explicit set $\mathcal{S}$ for which any minimal dendric shift (over some fixed alphabet) has an $\mathcal{S}$-adic representation, restricting ourselves to subfamilies of minimal dendric shifts allows to build examples. Furthermore, and in accordance to what was stated in the beginning of the previous subsection, if we consider all possible multi-cliques as the vertices of $\mathcal{G}^{L}(\mathcal{S})$ and $\mathcal{G}^{R}(\mathcal{S})$ (and not just the covering trees), any primitive $\mathcal{S}$-adic representation ${\boldsymbol{\sigma}}$ of $X$ will label, among others, the paths $(G^{L}(X_{\boldsymbol{\sigma}}^{(n)}))_{n\geq 0}$ and $(G^{R}(X_{\sigma}^{(n)}))_{n\geq 0}$ (in $\mathcal{G}^{L}(\mathcal{S})$ and $\mathcal{G}^{R}(\mathcal{S})$ respectively).
@@
As a consequence, it makes sense to look for an $S$-adic characterization of $F$ using infinite paths. It also implies that the elements of $F$ have $\mathcal{S}_{1\text{R}}$-adic representations where $\mathcal{S}_{1\text{R}}$ is the set of dendric return morphisms for a letter which has four right extensions. These morphisms can easily be found using the possible extension graphs for the empty word.
@@
The associated Rauzy graph of order 1 is
@@
and the return words for 00, corresponding to paths from and to the vertex 00, are given by 00, $01$, $02$ and $032$. Thus, the morphisms of $\mathcal{S}$ associated with this extension graph are the elements of $\beta\Sigma_{4}$ where $\beta$ is the morphism of Example 40, i.e. $\beta:0\mapsto 0,1\mapsto 01,2\mapsto 02,3\mapsto 032$, and $\Sigma_{4}$ is the set of permutations over $\{0,1,2,3\}$.
Up to a permutation of the alphabet, the possible extension graphs (with a unique right special letter) and their associated morphisms are given in Figure 6. Notice that, as we take the return morphisms for the only right special letter, there is, up to a permutation, only one morphism corresponding to each extension graph. The set $\mathcal{S}_{1\text{R}}$ is then finite and given by $\Sigma_{4}\{\alpha,\beta,\gamma,\delta\}\Sigma_{4}$, where $\alpha,\beta,\gamma,\delta$ are defined in Figure 6.
@@
We now build the graph that will give us the characterization. For the right side, we consider that the graph $\mathcal{G}^{R}(\mathcal{S}_{1\text{R}})$ is built with all the right valid triplets. However, as we have observed before, any $\mathcal{S}_{1\text{R}}$-adic representation ${\boldsymbol{\sigma}}$ of $X\in F$ labels the path $(G^{R}(X_{\boldsymbol{\sigma}}^{(n)}))_{n\geq 0}=(K_{4})_{n\geq 0}$. We can thus assume that the right side graph is reduced to the vertex $K_{4}$ without losing $\mathcal{S}_{1\text{R}}$-adic representations.
For the left side, we consider the usual graph $\mathcal{G}^{L}(\mathcal{S}_{1\text{R}})$ (with trees as vertices). When taking the product with the right side graph, we can use permutations and, since $K_{4}$ is symmetric, this gives us a graph with only two vertices $(G_{1},K_{4})$ and $(G_{2},K_{4})$ where $G_{1}$ and $G_{2}$ are represented in Figure 7.
@@
In the end, we obtain that each minimal shift space in $F$ has a primitive $\mathcal{S}_{1\text{R}}$-adic representation in the graph represented in Figure 8. Note that, to reduce the number of morphisms, we added a loop on $(G_{1},K_{4})$ (resp., $(G_{2},K_{4})$) labeled with the permutations that fix the graph $G_{1}$ (resp., $G_{2}$). We will then consider that $\mathcal{S}_{1\text{R}}$ also contains the permutations. To improve the readability of the graph, we also introduce the following notations: $\Sigma_{\{1,2,3\}}$ is the subset of permutations of $\Sigma_{4}$ fixing 00, $\pi_{ab}$ stands for the cyclic permutation $(a\ \ b)$ and we define the following morphisms
@@
Figure 8. A shift space over $\{0,1,2,3\}$ is minimal, dendric and has exactly one right special factor of each length if and only if it has a primitive $\mathcal{S}_{1\text{R}}$-adic representation labeling an infinite path in this graph.
The converse result is also true, i.e. if a shift space $X$ has a primitive $\mathcal{S}_{1\text{R}}$-adic representation ${\boldsymbol{\sigma}}$ labeling an infinite path in the graph of Figure 8, then it is a minimal dendric shift space over $\{0,1,2,3\}$ having exactly one right special factor of each length. Indeed, it is minimal dendric by the proof of Theorem 1. To conclude, we show that we can find infinitely many factors having four right extensions in $X$. This will imply that, for each length, there is a factor having four right extensions and it must be the unique right special factor since $X$ is dendric.
It follows from this observation: if $v$ has a left extension $a$ such that $av$ has four right extensions and if $\sigma\in\mathcal{S}_{1\text{R}}$ is a return morphism for $\ell$, then the extended image $\sigma(v)\ell$ has a left extension $b$ such that $b\sigma(v)\ell$ has four right extensions. Indeed, it suffices to take $b=\varphi^{L}_{\sigma,\varepsilon}(a)$ by Proposition 12 and by definition of the morphisms of $\mathcal{S}_{1\text{R}}$. For any $k>0$, starting with $\varepsilon$ in $X_{\boldsymbol{\sigma}}^{(k)}$, we can then find a word $u\in\mathcal{L}(X)$ of length at least $k$ having four right extensions.
@@ 7.4. Characterization of eventually dendric shifts
A minimal shift $X$ is eventually dendric if and only if it has a dendric derived shift, if and only if every derived shift with respect to a long enough factor of $X$ is dendric.
@@
A shift space is minimal eventually dendric if and only if it has a primitive $S$-adic representation $(\sigma_{n})_{n\geq 0}$ where $\sigma_{0}$ is a return morphism, $\sigma_{n}$ is a dendric return morphism for all $n\geq 1$ and $(\sigma_{n})_{n\geq 1}$ labels infinite paths in $\mathcal{G}^{L}(\mathcal{S})$ and $\mathcal{G}^{R}(\mathcal{S})$ where $\mathcal{S}=\{\sigma_{n}:n\geq 1\}$.
@@ 7.5. Decidability of eventual dendricity for minimal substitutive shifts
It is classical [15] that we can always suppose that $\sigma$ is non-erasing ($\sigma(a)\neq\varepsilon$ for every letter) and $\tau$ is a coding ($|\tau(a)|=1$ for every letter) so we always make such an assumption.
Observe that one usually considers the restricted version of morphic shifts consisting in shifts generated by a morphic word. The shift space generated by an infinite word $x\in\mathcal{A}^{\mathbb{N}}$ is the shift space $X_{x}=\{y\in\mathcal{A}^{\mathbb{Z}}\mid\mathcal{L}(y)\subset\mathcal{L}(x)\}$. Whenever $x$ is uniformly recurrent (that is, every finite word $u\in\mathcal{L}(x)$ occurs infinitely many times in $x$ and with bounded gaps), the shift space $X_{x}$ is minimal. An infinite word $x$ is morphic if there exist morphisms $\sigma:\mathcal{B}^{*}\to\mathcal{B}^{*}$ and $\tau:\mathcal{B}^{*}\to\mathcal{A}^{*}$ such that $\sigma$ is prolongable on some letter $b\in\mathcal{B}$ (that is, $\sigma(b)\in b\mathcal{B}^{*}$ and $\lim_{n\to+\infty}|\sigma^{n}(b)|=+\infty$) and $x=\tau(\sigma^{\omega}(b))=\lim_{n\to+\infty}\tau(\sigma^{n}(b^{\omega}))$. It is also decidable whether a morphic word is uniformly recurrent [23]. Though, the shift space $X_{x}$ could be minimal with $x$ not uniformly recurrent. Thus, deciding if $X(\sigma,\tau)$ is minimal is an open problem.
@@
If $X(\sigma,\tau)$ is finite, then it is generated by $x=u^{\omega}$, where $u$ is a finite word. It is direct to check that $x=\sigma^{\prime\omega}(b^{\prime})$, where $b^{\prime}$ is the first letter of $u$ and $\sigma(a)=u^{2}$ for every letter $a$.
@@
The set $B_{g}$ is non-empty by aperiodicity, so let $a\in B_{g}$. For every $n\in\mathbb{N}$, we set $\sigma^{n}(a)=u_{n}a_{n}v_{n}$, where $a_{n}\in B_{g}$ and $u_{n}\in B_{b}^{*}$. By the pigeonhole principle, we can find $k<\ell$ such that $a_{k}=a_{\ell}=b^{\prime}$. We thus have $\sigma^{\ell-k}(b^{\prime})=ub^{\prime}v$ for some word $u\in B_{b}^{*}$. Since $\sigma,\tau$ are non-erasing and since $X(\sigma,\tau)$ is aperiodic and minimal, we must have $u=\varepsilon$. In other words, $\sigma^{\prime}=\sigma^{\ell-k}$ is prolongable on $b^{\prime}$ and, by minimality, the word $\tau(\sigma^{\prime\omega}(b^{\prime}))$ generates $X(\sigma,\tau)$. ∎
@@
We now show that we can compute the dendricity threshold of $X_{x}$. As $X_{y}$ is dendric, if $G^{L}_{N}(X_{y})=G^{L}(X_{y})$ and $G^{R}_{N}(X_{y})=G^{R}(X_{y})$, then any word of $\mathcal{L}_{\geq N}(X_{y})$ is ordinary. Knowing $G^{L}(X_{y})$ and $G^{R}(X_{y})$, we can find such $N$ by iteratively computing the graphs $G^{L}_{n}(X_{y})$ and $G^{R}_{n}(X_{y})$. By Proposition 19, this gives an upper bound on the dendricity threshold of $X_{x}$. We can then compute its precise value by checking all small factors. ∎

### v3

@@ 7.2. Characterization via two graphs
Now assume that ${\boldsymbol{\sigma}}$ is primitive and labels a path $\left(G^{L}_{n}\right)_{n\geq 0}$ in $\mathcal{G}^{L}(\mathfrak{S})$ and a path $\left(G^{R}_{n}\right)_{n\geq 0}$ in $\mathcal{G}^{R}(\mathfrak{S})$. The shift $X_{\boldsymbol{\sigma}}$ is minimal by primitiveness of ${\boldsymbol{\sigma}}$ [21]. Let $u\in\mathcal{L}(X_{\boldsymbol{\sigma}})$ be a factor. If $u$ is a $\sigma_{0}$-initial factor, then it is dendric since $\sigma_{0}$ is dendric. Otherwise, iterating Proposition 8, there exist a unique $k>0$ and a unique $v\in\mathcal{L}(X_{\boldsymbol{\sigma}}^{(k)})$ $\sigma_{k}$-initial such that $u$ is an extended image of $v$ by $\sigma_{0}\dots\sigma_{k-1}$. Let $Y$ be an Arnoux-Rauzy shift and let $Z=\sigma_{k}\cdot Y$. Since $v$ is $\sigma_{k}$-initial, $\mathcal{E}_{Z}(v)=\mathcal{E}_{X_{\boldsymbol{\sigma}}^{(k)}}(v)$. As $Y$ is an Arnoux-Rauzy shift, the graph $G^{L}(Y)$ (resp., $G^{R}(Y)$) is the complete (simple and monochromatic) graph on $\mathcal{A}$, thus $G^{L}_{k+1}$ (resp., $G^{R}_{k+1}$) is a covering subgraph. By Lemma 51 applied $k+1$ times, we obtain that $\sigma_{0}\dots\sigma_{k-1}\cdot Z$ is dendric thus, starting with the extension graph $\mathcal{E}_{Z}(v)$, $v$ only has dendric extended images under $\sigma_{0}\dots\sigma_{k-1}$ and $u$ is dendric. ∎
@@
By construction, on an alphabet $\mathcal{A}$ of size $d$, the graphs $\mathcal{G}^{L}(\mathfrak{S})$ and $\mathcal{G}^{R}(\mathfrak{S})$ have $d^{d-2}$ vertices (see sequence A000272 on OEIS) each, meaning that if we take the product of these graphs, we obtain a characterization with a unique graph $\mathcal{G}(\mathfrak{S})$ with $d^{2(d-2)}$ vertices.
@@
However, it is possible to reduce the number of vertices with a clever use of permutations. Indeed, the vertices of $\mathcal{G}(\mathfrak{S})$ are pairs of labeled trees and some of these pairs are equivalent, in the sense that they are equal up to a permutation of the letters. Thus, by keeping only one pair per equivalence class and adding permutations to the morphisms of the set $\mathfrak{S}$, we obtain a smaller graph.
@@
7.3. An example: $\mathfrak{S}$-adic characterization of minimal dendric shifts with one right special factor of each length
While we do not have an easy description of an explicit set $\mathfrak{S}$ for which any minimal dendric shift (over some fixed alphabet) has an $\mathfrak{S}$-adic representation, restricting ourselves to subfamilies of minimal dendric shifts allows to build examples. Furthermore, and in accordance to what was stated in the beginning of the previous subsection, if we consider all possible multi-cliques as the vertices of $\mathcal{G}^{L}(\mathfrak{S})$ and $\mathcal{G}^{R}(\mathfrak{S})$ (and not just the covering trees), any primitive $\mathfrak{S}$-adic representation ${\boldsymbol{\sigma}}$ of $X$ will label, among others, the paths $(G^{L}(X_{\boldsymbol{\sigma}}^{(n)}))_{n\geq 0}$ and $(G^{R}(X_{\sigma}^{(n)}))_{n\geq 0}$ (in $\mathcal{G}^{L}(\mathfrak{S})$ and $\mathcal{G}^{R}(\mathfrak{S})$ respectively).
@@
As a consequence, it makes sense to look for an $S$-adic characterization of $F$ using infinite paths. It also implies that the elements of $F$ have $\mathfrak{S}_{1\text{R}}$-adic representations where $\mathfrak{S}_{1\text{R}}$ is the set of dendric return morphisms for a letter which has four right extensions. These morphisms can easily be found using the possible extension graphs for the empty word.
@@
The associated Rauzy graph of order 1, i.e., the graph where there is an edge from a letter $a$ to a letter $b$ whenever $ab\in\mathcal{L}(X)$, is the following
@@
and the return words for 00, corresponding to paths from and to the vertex 00, are given by 00, $01$, $02$ and $032$. Thus, the morphisms of $\mathfrak{S}$ associated with this extension graph are the elements of $\beta\Sigma_{4}$ where $\beta$ is the morphism of Example 40, i.e. $\beta:0\mapsto 0,1\mapsto 01,2\mapsto 02,3\mapsto 032$, and $\Sigma_{4}$ is the set of permutations over $\{0,1,2,3\}$.
Up to a permutation of the alphabet, the possible extension graphs (with a unique right special letter) and their associated morphisms are given in Figure 6. Note that, as we take the return morphisms for the only right special letter, there is, up to a permutation, only one morphism corresponding to each extension graph. The set $\mathfrak{S}_{1\text{R}}$ is then finite and given by $\Sigma_{4}\{\alpha,\beta,\gamma,\delta\}\Sigma_{4}$, where $\alpha,\beta,\gamma,\delta$ are defined in Figure 6.
@@
We now build the graph that will give us the characterization. For the right side, we consider that the graph $\mathcal{G}^{R}(\mathfrak{S}_{1\text{R}})$ is built with all the right valid triplets. However, as we have observed before, any $\mathfrak{S}_{1\text{R}}$-adic representation ${\boldsymbol{\sigma}}$ of $X\in F$ labels the path $(G^{R}(X_{\boldsymbol{\sigma}}^{(n)}))_{n\geq 0}=(K_{4})_{n\geq 0}$. We can thus assume that the right side graph is reduced to the vertex $K_{4}$ without losing $\mathfrak{S}_{1\text{R}}$-adic representations.
For the left side, we consider the usual graph $\mathcal{G}^{L}(\mathfrak{S}_{1\text{R}})$ (with trees as vertices). When taking the product with the right side graph, we can use permutations and, since $K_{4}$ is symmetric, this gives us a graph with only two vertices $(G_{1},K_{4})$ and $(G_{2},K_{4})$ where $G_{1}$ and $G_{2}$ are represented in Figure 7.
@@
In the end, we obtain that each minimal shift space in $F$ has a primitive $\mathfrak{S}_{1\text{R}}$-adic representation in the graph represented in Figure 8. Note that, to reduce the number of morphisms, we added a loop on $(G_{1},K_{4})$ (resp., $(G_{2},K_{4})$) labeled with the permutations that fix the graph $G_{1}$ (resp., $G_{2}$). We will then consider that $\mathfrak{S}_{1\text{R}}$ also contains the permutations. To improve the readability of the graph, we also introduce the following notations: $\Sigma_{\{1,2,3\}}$ is the subset of permutations of $\Sigma_{4}$ fixing 00, $\pi_{ab}$ stands for the cyclic permutation $(a\ \ b)$ and we define the following morphisms
@@
Figure 8. A shift space over $\{0,1,2,3\}$ is minimal, dendric and has exactly one right special factor of each length if and only if it has a primitive $\mathfrak{S}_{1\text{R}}$-adic representation labeling an infinite path in this graph.
The converse result is also true, i.e. if a shift space $X$ has a primitive $\mathfrak{S}_{1\text{R}}$-adic representation ${\boldsymbol{\sigma}}$ labeling an infinite path in the graph of Figure 8, then it is a minimal dendric shift space over $\{0,1,2,3\}$ having exactly one right special factor of each length. Indeed, it is minimal dendric by the proof of Theorem 1. To conclude, we show that we can find infinitely many factors having four right extensions in $X$. This will imply that, for each length, there is a factor having four right extensions and it must be the unique right special factor since $X$ is dendric.
It follows from this observation: if $v$ has a left extension $a$ such that $av$ has four right extensions and if $\sigma\in\mathfrak{S}_{1\text{R}}$ is a return morphism for $\ell$, then the extended image $\sigma(v)\ell$ has a left extension $b$ such that $b\sigma(v)\ell$ has four right extensions. Indeed, it suffices to take $b=\varphi^{L}_{\sigma,\varepsilon}(a)$ by Proposition 12 and by definition of the morphisms of $\mathfrak{S}_{1\text{R}}$. For any $k>0$, starting with $\varepsilon$ in $X_{\boldsymbol{\sigma}}^{(k)}$, we can then find a word $u\in\mathcal{L}(X)$ of length at least $k$ having four right extensions.
@@ 7.4. Characterization of eventually dendric shifts
A minimal shift $X$ is eventually dendric if and only if it has a dendric derived shift, or equivalently, if and only if every derived shift with respect to a long enough factor of $X$ is dendric.
@@
A shift space is minimal eventually dendric if and only if it has a primitive $S$-adic representation $(\sigma_{n})_{n\geq 0}$ where $\sigma_{0}$ is a return morphism, $\sigma_{n}$ is a dendric return morphism for all $n\geq 1$ and $(\sigma_{n})_{n\geq 1}$ labels infinite paths in $\mathcal{G}^{L}(\mathfrak{S})$ and $\mathcal{G}^{R}(\mathfrak{S})$ where $\mathfrak{S}=\{\sigma_{n}:n\geq 1\}$.
@@ 7.5. Decidability of eventual dendricity for minimal substitutive shifts
It is classical [15] that we can suppose that $\sigma$ is non-erasing ($\sigma(a)\neq\varepsilon$ for every letter) and $\tau$ is a coding ($|\tau(a)|=1$ for every letter) so we always make such an assumption.
Observe that one usually considers the restricted version of morphic shifts consisting in shifts generated by a morphic word. The shift space generated by an infinite word $x\in\mathcal{A}^{\mathbb{N}}$ is the shift space $X_{x}=\{y\in\mathcal{A}^{\mathbb{Z}}\mid\mathcal{L}(y)\subseteq\mathcal{L}(x)\}$. Whenever $x$ is uniformly recurrent (that is, every finite word $u\in\mathcal{L}(x)$ occurs infinitely many times in $x$ and with bounded gaps), the shift space $X_{x}$ is minimal. An infinite word $x$ is morphic if there exist morphisms $\sigma:\mathcal{B}^{*}\to\mathcal{B}^{*}$ and $\tau:\mathcal{B}^{*}\to\mathcal{A}^{*}$ such that $\sigma$ is prolongable on some letter $b\in\mathcal{B}$ (that is, $\sigma(b)\in b\mathcal{B}^{*}$ and $\lim_{n\to+\infty}|\sigma^{n}(b)|=+\infty$) and $x=\tau(\sigma^{\omega}(b))=\lim_{n\to+\infty}\tau(\sigma^{n}(b^{\omega}))$. It is also decidable whether a morphic word is uniformly recurrent [23]. Though, it is not a requirement for it to generate a minimal shift space, i.e., a non uniformly recurrent word $x$ could generate a minimal shift space. Thus, deciding if $X(\sigma,\tau)$ is minimal is an open problem.
@@
If $X(\sigma,\tau)$ is finite, then it is generated by $x=u^{\omega}$, where $u$ is a finite word. It is direct to check that $x=\sigma^{\prime\omega}(b^{\prime})$, where $b^{\prime}$ is the first letter of $u$ and $\sigma^{\prime}(a)=u^{2}$ for every letter $a$.
@@
The set $B_{g}$ is non-empty by aperiodicity, so let $a\in B_{g}$. For every $n\in\mathbb{N}$, we set $\sigma^{n}(a)=u_{n}a_{n}v_{n}$, where $a_{n}\in B_{g}$ and $u_{n}\in B_{b}^{*}$. By the pigeonhole principle, we can find $k<\ell$ such that $a_{k}=a_{\ell}=b^{\prime}$. We thus have $\sigma^{\ell-k}(b^{\prime})=ub^{\prime}v$ for some word $u\in B_{b}^{*}$, and $\sigma^{n(\ell-k)}(b^{\prime})$ starts with $\sigma^{(n-1)(\ell-k)}(u)\sigma^{(n-2)(\ell-k)}(u)\cdots\sigma^{\ell-k}(u)u$ for all $n$. As $u\in B_{b}^{*}$, the sequence $(\sigma^{n(\ell-k)}(u))_{n\geq 0}$ is eventually periodic. Hence, as $\sigma$ and $\tau$ are non-erasing, $u=\varepsilon$ otherwise this contradicts the aperiodicity of $X(\sigma,\tau)$. In other words, $\sigma^{\prime}=\sigma^{\ell-k}$ is prolongable on $b^{\prime}$ and, by minimality, the word $\tau(\sigma^{\prime\omega}(b^{\prime}))$ generates $X(\sigma,\tau)$. ∎
@@
We now show that we can compute the dendricity threshold of $X_{x}$. As $X_{y}$ is dendric, if $G^{L}_{N}(X_{y})=G^{L}(X_{y})$ and $G^{R}_{N}(X_{y})=G^{R}(X_{y})$, then any word of $\mathcal{L}_{\geq N}(X_{y})$ is ordinary. Knowing $G^{L}(X_{y})$ and $G^{R}(X_{y})$, we can find such $N$ by iteratively computing the graphs $G^{L}_{n}(X_{y})$ and $G^{R}_{n}(X_{y})$ and comparing them to $G^{L}(X_{y})$ and $G^{R}(X_{y})$. By Proposition 19, this gives an upper bound on the dendricity threshold of $X_{x}$. We can then compute its precise value by checking all small factors. ∎